# Bones of Contention: Exploring Query-Efficient Attacks Against Skeleton Recognition Systems

### v1

@@ Abstract
Skeleton action recognition models have secured more attention than video-based ones in various applications due to privacy preservation and lower storage requirements. Skeleton data are typically transmitted to cloud servers for action recognition, with results returned to clients via Apps/APIs. However, the vulnerability of skeletal models against adversarial perturbations gradually reveals the unreliability of these systems. Existing black-box attacks all operate in a decision-based manner, resulting in numerous queries that hinder efficiency and feasibility in real-world applications. Moreover, all attacks off the shelf focus on only restricted perturbations, while ignoring model weaknesses when encountered with non-semantic perturbations. In this paper, we propose two query-effIcient Skeletal Adversarial AttaCks, ISAAC-K and ISAAC-N. As a black-box attack, ISAAC-K utilizes Grad-CAM in a surrogate model to extract key joints where minor sparse perturbations are then added to fool the classifier. To guarantee natural adversarial motions, we introduce constraints of both bone length and temporal consistency. ISAAC-K finds stronger adversarial examples on $\ell_{\infty}$ norm, which can encompass those on other norms. Exhaustive experiments substantiate that ISAAC-K can uplift the attack efficiency of the perturbations under 10 skeletal models. Additionally, as a byproduct, ISAAC-N fools the classifier by replacing skeletons unrelated to the action. We surprisingly find that skeletal models are vulnerable to large perturbations where the part-wise non-semantic joints are just replaced, leading to a query-free no-box attack without any prior knowledge. Based on that, four adaptive defenses are eventually proposed to improve the robustness of skeleton recognition models.
Index Terms:
Adversarial attacks, skeleton recognition, query-efficient attacks.
@@ I Introduction
As an important branch of computer vision, action recognition has been widely used in applications encompassing aged care [1], human-computer interaction [2], video surveillance [3], intelligent transportation [4], to name a few. In Metaverse systems, most virtual reality devices are equipped with joint sensors to obtain human posture in human-computer interaction [5, 6]. Compared to traditional RGB video action recognition methods [7, 8, 9], skeleton-based action recognition methods can greatly reduce the amount of computation and storage space, avoid the impact of illumination and occlusion [10, 11], and, more importantly, prevent privacy disclosure to some extent [12]. Skeleton data can be easily obtained using joint sensors [5, 6] or devices like Kinect cameras [13], positioning it as a new competitive modality. However, the robustness of skeletal models is not fully unveiled. Once compromised, these models may pose a grave threat to mission-critical environments and cause property loss and casualties. Hence, research into the robustness of skeleton recognition systems assumes paramount significance for ensuring secure and reliable applications.
@@
The security scenario of this paper is illustrated in Figure 1. Most recognition systems delegate data storage and model training/inference to outsourcing companies due to unbearable burden. Concretely, after the skeleton data are obtained by sensors, they are uploaded to the cloud server for action recognition. Finally, the output is sent to the client through Apps/APIs, e.g., Kinect SDK and Action Rekonition. However, the reliance on third-party outsourcing services raises significant concerns about data privacy and security. We consider two ways of attacks in this process. 1) Man-in-the-middle manipulation: the attacker can intercept the skeleton received on the cloud server and replace it with elaborately designed adversarial examples. A strong motivation of this attack is to conceal urgent events captured by surveillance cameras, such as critical situations in aged care or illicit activities, by hindering these incidents from being easily detected or flagged. This attack also illustrates that the security of information still remains a significant concern, although the data have been desensitized. 2) Maneuvering manipulation: the person whose pose is tracked (not necessarily the client, e.g., in video surveillance) can act as an attacker. For example, the perpetrators are aware that their behavior is monitored, but they can make an attempt to fool the recognition model to some extent by frequently changing non-semantic poses since we find that by simply replacing lower body with various postures, existing skeletal models tend to output different labels for the same upper-body action such as drinking water (see more details in Section V-E).
@@
TABLE I: A comparison of adversarial attacks against skeleton action recognition models.
@@
Kumar [14]
Liu [15]
Zheng [16]
Wang [17]
Tanaka [18]
Huang [19]
BASAR [20]
QESAR [21]
Lu [22]
TASAR [23]
ISAAC-K (ours) 1
@@
Figure 1: Attacking Scenario. A typical skeleton recognition system includes a local service (extracting skeleton from video and receiving system output through Apps/APIs) and a cloud service (storing the skeleton on a cloud server and inputting it into the recognition model). The attack can occur on both the cloud server and original videos.
@@
The above scenarios drive us to search for adversarial perturbations for skeleton data. Despite a line of existing skeleton action recognition attacks (as listed in Table I), they are encountered with three main limitations. 1) Most attacks are under a white-box setting [14, 15, 16, 17, 18, 19], which requires a wealth of prior knowledge of the victim model. However, in actual scenarios, most commercial models do not disclose their model architecture, so accessibility to the model is impractical, which necessitates the emergence of black-box attacks where attackers can only obtain the model output. Despite few gray-box attacks that rest upon the transferability of skeletal models, their performance is quite weak and susceptible to the selected surrogate model [23, 24]. 2) Existing black-box attacks are lacking in attack efficiency. BASAR [20] and QESAR [21] are both decision-based black-box attacks, which require a large number of queries due to the search process. Moreover, they perturb all joints, while the perturbations on some inconsequential joints are not necessary, which significantly reduces the attack efficiency. In practice, attack efficiency counts, since large queries imply high attack costs. A later study [19] considers adding sparse perturbations to skeletons by extracting key joints with dynamic attention, but the attack is carried out under the white-box setting, which greatly overestimates the ability of the attacker. Note that squarely transferring white-box methods to black-box attacks poses challenges. This is primarily due to two reasons. First, unlike images, the dimensions of skeleton data are much lower; thus, diminutive perturbations at any joint may cause optimization to fall into local optima. Additionally, the lack of gradient information can easily lead to an impasse in the search for perturbation directions, ultimately leading to the inability to converge. 3) All off-the-shelf attacks consider restricted attacks that only add minuscule perturbations to the input motion. We regard this as reasonable, but it is not the only way to add perturbations since the non-semantic joints have no direct connection with the classification label. For a sufficiently robust classifier, the non-semantic joints added with large perturbations are not expected to affect the classification results as long as the whole skeleton conforms to the human manifold. Taking an action classified as “drinking water” as an example, no matter how the lower body changes (e.g., sitting, crouching, or kneeling), an expected classifier should always classify it as “drinking water”, even if the skeleton sample has not appeared in the training set.
@@
To close the gaps above, we first propose a query-effIcient Skeletal Adversarial AttaCk based on Key joints, called ISAAC-K. We only add perturbations to key joints, while the naturalness of adversarial motions is maintained by bone length (intra) and temporal consistency (inter) constraints. Afterwards, the skeletons are optimized by an improved version of RayS [25] to enhance attack efficiency. Experiments show that ISAAC-K performs better than the existing open-source State-Of-The-Art (SOTA) attack, BASAR, and can bypass existing adversarial defenses. For non-semantic joints, following the semantic invariant perturbations considered in the image/video domain [26, 27], we propose ISAAC-N and show that, even in a no-box setting (the attacker cannot even query the model), by making reasonable changes to the non-semantic joints, the model output can be misguided without affecting the human perception of skeleton actions. These perturbations, which surpass metrical restrictions, are marked as unrestricted perturbations [28], significantly enhancing their attacking performance and resistance to adaptive defenses. The proposed ISAAC-K and ISAAC-N are customized for man-in-the-middle manipulation and maneuvering manipulation delineated above, respectively.
@@
  • We propose a novel query-efficient black-box skeletal adversarial attack, ISAAC-K. We use Grad-CAM [29] to extract key joints from a local surrogate model and search stronger adversarial examples on the $\ell_{\infty}$ ball, as it is the superset of other $\ell_{p}$ balls under the same perturbation radius.
@@
  • As a byproduct, ISAAC-N show that even when the attacker cannot query skeletal models, they are surprisingly vulnerable to semantic invariant perturbations such as part-wise replacement of non-semantic joints. As responses, we propose four adaptive defenses to improve the robustness of skeletal models.
@@ II Related Work
Skeleton Attacks. Despite a plethora of action recognition methods based on skeletons, research on adversarial attacks against such models is still in its infancy. Most attacks against skeleton action recognition models are under the white-box setting, where the attacker can access the structure and parameters of the victim model. An early study [14] extracts key frames and key joints through feature information for multi-modal data including RGB videos, depth videos and skeleton data. Recent skeleton attacks preserve the imperceptibility of the adversarial skeleton by restricting either the bone length [18, 15], rotation angle [16], velocity [16], or joint accelerations [15, 17]. However, such minor changes may cause off-manifold skeletons that violate the basic requirements of human dynamics. To handle this, a black-box attack called BASAR [20, 30] is proposed to move the perturbed skeleton to the manifold. The random search-based attack, however, causes a large number of queries and low attack efficiency, which are not mentioned in the paper. Kang et al. [21] claim that they reduce query to some extent by introducing bias sampling, but the accuracy of gradient estimation is affected by the sampling number, and the constraints are not clarified in detail. Huang et al. [19] uses dynamic attention to select joints that are highly dynamic, but under a white-box setting. Moreover, the joints extracted by attention mechanism [14, 19] are not consistent with key joints closely related to classification, and the latter plays a greater role in affecting the model outputs. Lu et al. [22] propose a hard no-box attack in which the attacker cannot obtain the victim model and labels. However, the attack success rate remains low, which makes this difficult attack setting not effective. A concurrent work [23] explores the transferability of skeletal adversarial examples in the gray-box setting, but suffers from low performance and sensitivity to surrogate models.
@@
Skeleton Defenses. There is also a lack of research on skeleton defenses, leaving space for adversaries to attack surreptitiously without being detected. Zheng et al. [16] design an effective plug-in defense by adding Gaussian noise and temporal filtering. Wang et al. [31] propose a black-box defense, BEAT, which discriminates adversarial motions from the view of joint Bayesian together with a perturbation restriction strategy based on natural motion manifolds. BEAT can successfully resist BASAR and outstrip competitors such as Zheng et al. [16] and adversarial training (AT) [32] by improving the classifier’s robustness and reducing the memory footprint. Hu et al. [33] devise a graph convolution network for multilevel skeleton forgery detection to recognize the tampering anomaly, but it is not applicable to adversarial perturbations. Recently, Tanaka et al. [34] tailor a Fourier-based AT method for $\ell_{2}$-PGD attack to improve model robustness.
@@ III Preliminary
Problem Statement. We denote the skeleton action recognition model as $f\left({\mathbf{x}}\right):\mathbf{x}\to y$, which takes a motion $\mathbf{x}\in{\mathbb{R}^{T\times N\times D}}$ as input, and outputs the predicted label $y$, where $T$, $N$, and $D$ represent the frame number, degree of freedoms (usually positions or angles of joints) and joint dimension, respectively. The attacker then regards the attack as an optimization problem by solving:
$\begin{array}[]{c}\mathop{\arg\min}\limits_{{\mathbf{x}_{adv}}}L\left({\mathbf{x},{\mathbf{x}_{adv}}}\right),\\ \mathrm{s.t.}\quad{\mathbf{x}_{adv}}\in\left[{0,1}\right],\ \ {f\left({{\mathbf{x}_{adv}}}\right)\neq y},\end{array}$ (1)
where $L$ denotes the distance function which ensures that the adversarial motion $\mathbf{x}_{adv}$ is close to the original motion $\mathbf{x}$. Note that it is common practice to rescale the data to $\left[{0,1}\right]$ before feeding it into models. The adversarial motion is expected to be predicted as any class excluding the ground truth class $y$ in untargeted attacks, while the pre-determined class in targeted attacks. Solving the above non-linear problem is difficult and requires substitutional means. The previous work, BASAR [20], utilizes a numerical solution and moves the adversarial motion to the human manifold $\mathcal{M}$ to maintain naturalness. In our work, we modify the above optimization function and adapt it to two different sparse attacks, ISAAC-K and ISAAC-N. ISAAC-K denotes the sparse attack in which perturbations are only added to the key joints, the framework of which is depicted in Figure 2. The key joints are first extracted by a local surrogate model, after which the perturbations are optimized iteratively under the constraints of bone length and temporal consistency. As a derivative of ISAAC-K, ISAAC-N represents the sparse attack in which perturbations are added to non-semantic joints, which will be introduced in Section IV-C.
@@
Threat Model. We assume that the attacker mounts the attack under black-box settings, where there is no access to the architecture and parameters of the victim model. More strictly, we consider hard-label attacks (consistent with BASAR [20]), which indicates that the attacker can only obtain the Top-1 label when querying the model. We also set an upper query limit for query, which emulates the limit imposed by commercial action recognition systems in the real world. Besides, in ISAAC-N, the attacker can replace the non-semantic joints by uploading the specific action taken by himself or from other online resources, instead of selecting from the dataset. This enables action customization for users and supports specified skeleton attacks.

### v2

@@ Abstract
Skeleton action recognition models have secured more attention than video-based ones in various applications due to privacy preservation and lower storage requirements. Skeleton data are typically transmitted to cloud servers for action recognition, with results returned to clients via Apps/APIs. However, the vulnerability of skeletal models against adversarial perturbations gradually reveals the unreliability of these systems. Existing black-box attacks all operate in a decision-based manner, resulting in numerous queries that hinder efficiency and feasibility in real-world applications. Moreover, all attacks off the shelf focus on only restricted perturbations, while ignoring model weaknesses when encountered with non-semantic perturbations. In this paper, we propose two query-effIcient Skeletal Adversarial AttaCks, ISAAC-K and ISAAC-N. As a black-box attack, ISAAC-K utilizes Grad-CAM in a surrogate model to extract key joints where minor sparse perturbations are then added to fool the classifier. To guarantee natural adversarial motions, we introduce constraints of both bone length and temporal consistency. ISAAC-K finds stronger adversarial examples on the $\ell_{\infty}$ norm, which can encompass those on other norms. Exhaustive experiments substantiate that ISAAC-K can uplift the attack efficiency of the perturbations under 10 skeletal models. Additionally, as a byproduct, ISAAC-N fools the classifier by replacing skeletons unrelated to the action. We surprisingly find that skeletal models are vulnerable to large perturbations where the part-wise non-semantic joints are just replaced, leading to a query-free no-box attack without any prior knowledge. Based on that, four adaptive defenses are eventually proposed to improve the robustness of skeleton recognition models.
@@ I Introduction
As an important branch of computer vision, action recognition has been widely used in applications encompassing aged care [jang2020etri], human-computer interaction [ann2014human], video surveillance [han2018going], intelligent transportation [xu2021action], to name a few. In Metaverse systems, most virtual reality devices are equipped with joint sensors to obtain human posture in human-computer interaction [jiang2022avatarposer, dai2023sparse]. Compared to traditional RGB video action recognition methods [donahue2015long-term, wang2016TSN, hara2018R3D], skeleton-based action recognition methods can greatly reduce the amount of computation and storage space, avoid the impact of illumination and occlusion [yue2022action, zheng2024towards], and, more importantly, prevent privacy disclosure to some extent [moon2023anonymization]. Skeleton data can be easily obtained using joint sensors [jiang2022avatarposer, dai2023sparse] or devices like Kinect cameras [zhang2012microsoft], positioning it as a new competitive modality. However, the robustness of skeletal models is not fully unveiled. Unlike pixel-level manipulations in images, perturbations in joint positions can preserve natural appearance yet still lead to severe semantic misclassification. Once compromised, such as failing to detect a fall incident in aged care, or misinterpreting suspicious behavior in a security context, these models may pose a grave threat to mission-critical environments and cause property loss and casualties. Hence, research into the robustness of skeleton recognition systems assumes paramount significance for ensuring secure and reliable applications.
@@
The security scenario of this paper is illustrated in Fig. 1(a). Most recognition systems delegate data storage and model training/inference to outsourcing companies due to unbearable burden. Concretely, after the skeleton data are obtained by sensors, they are uploaded to the cloud server for action recognition. Finally, the output is sent to the client through Apps/APIs, e.g., Kinect SDK and Amazon Rekognition. However, the reliance on third-party outsourcing services raises significant concerns about data privacy and security. We consider two ways of attacks in this process. 1) Man-in-the-middle manipulation: the attacker can intercept the skeleton received on the cloud server and replace it with elaborately designed adversarial examples. A strong motivation of this attack is to conceal urgent events captured by surveillance cameras, such as critical situations in aged care or illicit activities, by hindering these incidents from being easily detected or flagged. 2) Maneuvering manipulation: the person whose pose is tracked (not necessarily the client) can act as an attacker. For example, the perpetrators are aware that their behavior is monitored, but they can make an attempt to fool the recognition model to some extent by frequently changing non-semantic poses since we find that by simply replacing lower body with various postures, existing skeletal models tend to output different labels for the same upper-body action such as drinking water (see more details in Section V-E).
Figure 1: (a) Attacking Scenario. A typical skeleton recognition system includes a local service (extracting skeleton from video and receiving system output through Apps/APIs) and a cloud service (storing the skeleton on a cloud server and inputting it into the recognition model). The attack can occur on both the cloud server and original videos. (b) Practical examples of misclassification.
To concretely demonstrate the feasibility and potential harm of such attacks, Fig. 1(b) shows practical cases where adversarial perturbations cause critical misclassifications in workplace monitoring and elderly care. For instance, actions such as sitting on the floor or falling are incorrectly recognized as benign gestures such as “put the palms together” or “playing with phone/tablet”, potentially leading to undetected misconduct or missed emergency alerts. These examples reveal how adversarial skeleton manipulation, even under constrained visual cues, can undermine the trustworthiness of action recognition systems deployed in safety-critical environments.
@@
TABLE I: Comparison of adversarial attacks against skeleton action recognition models.
@@
Kumar [kumar2020finding]
Liu [liu2020adversarial]
Zheng [zheng2020towards]
Wang [wang2021understanding]
Tanaka [tanaka2022adversarial]
Huang [huang2022sparse]
BASAR [diao2021basar]
QESAR [kang2023qesar]
Lu [lu2023hard]
TASAR [diao2024tasar]
ISAAC-K (ours)
1
@@
The above scenarios drive us to search for adversarial perturbations for skeleton data. Despite a line of existing skeleton action recognition attacks (as listed in Table I), they are encountered with three main limitations. 1) Most attacks are under a white-box setting [kumar2020finding, liu2020adversarial, zheng2020towards, wang2021understanding, tanaka2022adversarial, huang2022sparse], which requires a wealth of prior knowledge of the victim model. However, in actual scenarios, most commercial models do not disclose their model architecture, so accessibility to the model is impractical, which necessitates the emergence of black-box attacks where attackers can only obtain the model output. Despite few gray-box attacks that rest upon the transferability of skeletal models, their performance is quite weak and susceptible to the selected surrogate model [diao2024tasar, sun2024invisibility]. 2) Existing black-box attacks are lacking in attack efficiency. BASAR [diao2021basar] and QESAR [kang2023qesar] are both decision-based black-box attacks, which require a large number of queries due to the search process. Moreover, they perturb all joints, while the perturbations on some inconsequential joints are not necessary, which significantly reduces the attack efficiency. In practice, attack efficiency counts, since large queries imply high attack costs. A later study [huang2022sparse] considers adding sparse perturbations to skeletons by extracting key joints with dynamic attention, but the attack is carried out under the white-box setting, which greatly overestimates the ability of the attacker. Note that squarely transferring white-box methods to black-box attacks poses challenges. This is primarily due to two reasons. First, unlike images, the dimensions of skeleton data are much lower; thus, diminutive perturbations at any joint may cause optimization to fall into local optima. Additionally, the lack of gradient information can easily lead to an impasse in the search for perturbation directions, ultimately leading to the inability to converge. 3) All off-the-shelf attacks consider restricted attacks that only add minuscule perturbations to the input motion. We regard this as reasonable, but it is not the only way to add perturbations since the non-semantic joints have no direct connection with the classification label. For a sufficiently robust classifier, the non-semantic joints added with large perturbations are not expected to affect the classification results as long as the whole skeleton conforms to the human manifold. Taking an action classified as “drinking water” as an example, no matter how the lower body changes (e.g., sitting, crouching, or kneeling), an expected classifier should always classify it as “drinking water”, even if the skeleton sample has not appeared in the training set.
@@
To close the gaps above, we first propose a query-effIcient Skeletal Adversarial AttaCk based on Key joints, called ISAAC-K. We only add perturbations to key joints, while the naturalness of adversarial motions is maintained by two domain-specific constraints. The spatial bone length constraint maintains kinematic plausibility by recursively enforcing bone length consistency across child joints and parent joints, while the temporal consistency constraint ensures perturbation smoothness across adjacent frames by dynamically correcting perturbation directions based on temporal alignment. Afterwards, the skeletons are optimized by an improved version of RayS [chen2020rays], which substantially reduces the effective search complexity and enables faster optimization. Experiments show that ISAAC-K performs better than the existing open-source State-Of-The-Art (SOTA) attack, BASAR, and can bypass existing adversarial defenses. For non-semantic joints, following the semantic invariant perturbations considered in the image/video domain [shamsabadi2020colorfool, cao2023stylefool], we propose ISAAC-N and show that, even in a no-box setting (the attacker cannot even query the model), by making reasonable changes to the non-semantic joints, the model output can be misguided without affecting the human perception of skeleton actions. This part-wise substitution approach deviates from conventional perturbation-based frameworks and exposes a new form of vulnerability: the model’s over-reliance on part of the body. These perturbations, which surpass metrical restrictions, are marked as unrestricted perturbations [bhattad2020unrestricted], significantly enhancing their attacking performance and resistance to adaptive defenses. The proposed ISAAC-K and ISAAC-N are customized for man-in-the-middle manipulation and maneuvering manipulation delineated above, respectively. They also help expose structural vulnerabilities and provide valuable insights to enhance model robustness for model developers.
@@
  • We propose a novel query-efficient black-box skeletal adversarial attack, ISAAC-K. We use Grad-CAM [selvaraju2017gradcam] to extract key joints from a local surrogate model and search for stronger adversarial examples on the $\ell_{\infty}$ ball, as it is the superset of other $\ell_{p}$ balls under the same perturbation radius.
@@
  • As a byproduct, ISAAC-N reveals that even when the attacker cannot query skeletal models, they are surprisingly vulnerable to semantic invariant perturbations such as part-wise replacement of non-semantic joints. As responses, we propose four adaptive defenses to improve the robustness of skeletal models.
@@ II Related Work
Skeleton Attacks. Despite a plethora of action recognition methods based on skeletons [shi2020msaagcn, yan2018stgcn, xu2023lagcn, do2024skateformer], research on adversarial attacks against such models is still in its infancy. Most attacks against skeleton action recognition models are under the white-box setting, where the attacker can access the structure and parameters of the victim model. An early study [kumar2020finding] extracts key frames and key joints through feature information for multi-modal data including RGB videos, depth videos and skeleton data. Recent skeleton attacks preserve the imperceptibility of the adversarial skeleton by restricting either the bone length [tanaka2022adversarial, liu2020adversarial], rotation angle [zheng2020towards], velocity [zheng2020towards], or joint accelerations [liu2020adversarial, wang2021understanding]. However, such minor changes may cause off-manifold skeletons that violate the basic requirements of human dynamics. To handle this, a black-box attack called BASAR [diao2021basar] is proposed to move the perturbed skeleton to the manifold. This random search-based attack, however, causes a large number of queries and low attack efficiency, which are not mentioned in the paper. Kang et al. [kang2023qesar] claim that they reduce query to some extent by introducing bias sampling, but the accuracy of gradient estimation is affected by the sampling number, and the constraints are not clarified in detail. Huang et al. [huang2022sparse] uses dynamic attention to select joints that are highly dynamic, but under a white-box setting. Moreover, the joints extracted by attention mechanism [kumar2020finding, huang2022sparse] are not consistent with key joints closely related to classification, and the latter plays a greater role in affecting model outputs. Lu et al. [lu2023hard] propose a hard no-box attack in which the attacker cannot obtain the victim model and labels. However, the attack success rate remains low, which makes this difficult attack setting not effective. A recent work [diao2024tasar] explores the transferability of skeletal adversarial examples in the gray-box setting, but suffers from low performance and sensitivity to surrogate models.
@@
Skeleton Defenses. There is also a lack of research on skeleton defenses, leaving space for adversaries to attack surreptitiously without being detected. Zheng et al. [zheng2020towards] design an effective plug-in defense by adding Gaussian noise and temporal filtering. Wang et al. [wang2023defending] propose a black-box defense, BEAT, which discriminates adversarial motions from the view of joint Bayesian together with a perturbation restriction strategy based on natural motion manifolds. BEAT can successfully resist BASAR and outstrip competitors such as Zheng et al. [zheng2020towards] and adversarial training (AT) [madry2018towards] by improving the classifier’s robustness and reducing the memory footprint. Hu et al. [hu2023self] devise a graph convolution network for multilevel skeleton forgery detection to recognize anomalies, but it is not applicable to adversarial perturbations. Recently, Tanaka et al. [tanaka2024fourier] tailor a Fourier-based AT method for $\ell_{2}$-PGD attack to improve model robustness.
@@ III Preliminary
Problem Statement. We denote the skeleton action recognition model as $f\left({\mathbf{x}}\right):\mathbf{x}\to y$, which takes a motion $\mathbf{x}\in{\mathbb{R}^{T\times N\times D}}$ as input, and outputs the predicted label $y$, where $T$, NN, and $D$ represent the frame number, degree of freedoms (usually positions or angles of joints) and joint dimension, respectively. The attacker then regards the attack as an optimization problem by solving:
$\small\begin{array}[]{c}\mathop{\arg\min}\limits_{{\mathbf{x}_{adv}}}\hat{L}\left({\mathbf{x},{\mathbf{x}_{adv}}}\right),\qquad\\ \mathrm{s.t.}\qquad{\mathbf{x}_{adv}}\in\left[{0,1}\right],\ \ {f\left({{\mathbf{x}_{adv}}}\right)\neq y},\end{array}$ (1)
where $\hat{L}$ denotes the distance function which ensures that the adversarial motion $\mathbf{x}_{adv}$ is close to the original motion $\mathbf{x}$. Note that it is common practice to rescale the data to $\left[{0,1}\right]$ before feeding it into models. The adversarial motion is expected to be predicted as any class excluding the ground truth class $y$ in untargeted attacks, while the pre-determined class in targeted attacks. Solving the above non-linear problem is difficult and requires substitutional means. The previous work, BASAR [diao2021basar], utilizes a numerical solution and moves the adversarial motion to the human manifold $\mathcal{M}$ to maintain naturalness. In our work, we modify the above optimization function and adapt it to two different sparse attacks, ISAAC-K and ISAAC-N. ISAAC-K denotes the sparse attack in which perturbations are only added to the key joints, the framework of which is depicted in Fig. 2. The key joints are first extracted by a local surrogate model, after which the perturbations are optimized iteratively under the constraints of bone length and temporal consistency. As a derivative of ISAAC-K, ISAAC-N represents the sparse attack in which perturbations are added to non-semantic joints, which will be introduced in Section IV-C.
@@
Threat Model. We assume that the attacker mounts the attack under black-box settings, where there is no access to the architecture and parameters of the victim model. More strictly, we consider hard-label attacks (consistent with BASAR [diao2021basar]), which indicates that the attacker can only obtain the Top-1 label when querying the model. We also set an upper query limit, which emulates the constraint imposed by commercial action recognition systems in the real world. Besides, in ISAAC-N, the attacker can replace the non-semantic joints by uploading the specific action taken by himself or from other online resources, instead of selecting from the dataset. This enables action customization for users and supports specified skeleton attacks.